OpenStudy (anonymous):

How do you solve this: 4(2x-9)=3(6x+5)?

hartnn (hartnn):

use distributive property first: \[a(b+c)=ab+ac\] and tell me what u get for 4(2x-9)

OpenStudy (anonymous):

Hold on

OpenStudy (jiteshmeghwal9):

first distribute both sides 8x-36=18x+15 now take variables @ one side & constants @ one side 8x-18x=15+36 now try to solve for 'x'. may be ! u can do it :)